User name X. Jiang
Log entry time 00:18:29 on February 12,1999
Entry number 13815
keyword=scintillator inefficiency, after hv change.
Looked at scintillator inefficiency again , after the HV adjustment.
Run 1947. The T4/T3 ratio has dropped significantly from about 11% to about 8%.
The following is the adc spectra with good tdc cut (upper figure) and without good tdc cut (lower figure) for s2r2. Looks like the problem has been fixed for now.